# Representative Action Selection for Large Action-Space Meta-Bandits

### v1

@@ 1 Introduction
Different from prior approaches in multi-armed bandits (MAB) (dani2008stochastic, ) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits. This problem can be seen as a large-scale combinatorial optimization under uncertainty, with applications where decisions involve a vast number of possibilities but are constrained by computational or time limitations for evaluating all options, e.g., inventory management, online recommendations.
Consider the following MAB setting: In a bandit, if a decision-maker plays an action $a\in\mathcal{A}_{full}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$. We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle,$ where this bandit instance $\theta\in\mathbb{R}^{n}$ follows a multivariate Gaussian distribution. Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ forms a Gaussian process (vershynin2018high, ). Our underlying setting aligns with the one considered in contextual bandit (dani2008stochastic, ; bietti2021contextual, ); however, the key difference lies in the objective.
Now consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{full}$. For a given bandit instance $\theta$, if the best action over the full space lies within $\mathcal{A}$, the decision-maker benefits from a reduced set of suboptimal actions to explore. Conversely, if the best action lies outside $\mathcal{A}$, regret arises from being unable to select this best action. Thus, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
$\textnormal{Regret}:=\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$ (1)
@@
This objective is motivated by practical considerations. MAB regret in finite-horizon settings (auer2002using, ; auer2002nonstochastic, ) typically scales with the number of available actions. However, if a subset $\mathcal{A}$ is carefully chosen, the resulting MAB regret can be lower than when using the full action space, due to a reduced set of suboptimal actions. This is especially beneficial when there are a very large number of actions, as even the initialization phase can be computationally expensive. This improvement arises from a natural connection between the regret in Equation (1) and MAB regret, as discussed in Supplementary A.
The remainder of this section is literature review. The subsequent sections present our main contributions and are organized as follows:
  • Section 2: Subset Selection Framework and Algorithm. We introduce a framework that models expected outcome using a Gaussian process to capture correlations across similar actions and bandit instances. Based on this structure, we construct geometric $\epsilon$-nets and refine them using an importance-based selection mechanism. This mechanism is formalized as a variant of the $\epsilon$-net algorithm (haussler1986epsilon, ), tailored to the bandit setting.
  • Section 3: Regret Analysis. We provide theoretical guarantees, including upper and lower bounds on the expected regret that reflect the geometric and importance structure of the action space.
  • Section 4: Examples and Empirical Validation. We extend our analysis to settings where outcome (payoff) functions are sampled from a Gaussian process with a kernel, and empirically compare the performance to that of Thompson Sampling and Upper Confidence Bound.
The code used for the experiments is available at https://github.com/Quan-Zhou/MAB.
@@ 1.1 Related Works
Multi-Armed Bandits The classic MAB problem (lai1985asymptotically, ; auer2002finite, ) is defined by a set of arms, each deliver outcomes that are independently drawn from a fixed and unknown distribution. The decision-maker sequentially selects an arm, observes its outcome, and aims to maximize cumulative outcomes over time. The policy is to choose the next arm based on the sequence of past plays and obtained outcomes. Popular policies include the Upper Confidence Bound (auer2002finite, ), Thompson Sampling (agrawal2012analysis, ), and EXP3 (auer2002nonstochastic, ).
Contextual Bandits Unlike classic MAB, which assumes action outcomes are independent, contextual bandits (auer2002using, ) incorporate observed features (contexts) at each round. The decision-maker selects actions based on context, aiming to learn a policy that maps contexts to high-reward actions. Linear models such as LinUCB (chu2011contextual, ) and linear Thompson sampling (agrawal2013thompson, ) assume linear reward functions, enabling efficient learning. Early work includes LinRel (auer2002using, ), while later efforts extended to nonlinear models (filippi2010parametric, ; jun2017scalable, ). Nonlinear payoff structures have also been explored using Gaussian processes (srinivas2009gaussian, ) and RKHS methods (valko2013finite, ; chowdhury2017kernelized, ). Recent studies examine richer structures: budget-constrained linear contextual bandits (agrawal2016linear, ), reward feedback across contexts (balseiro2019contextual, ), and adversarially corrupted outcomes (he2022nearly, ). Empirical evaluations of common algorithms appear in bietti2021contextual .
Online Combinatorial Optimization Several works have extended stochastic linear optimization (dani2008stochastic, ; rusmevichientong2010linearly, ) to combinatorial settings. This line of work applies the Follow-the-Perturbed-Leader (FPL) algorithm (hannan1957approximation, ; kalai2005efficient, ), which introduces random perturbations to the estimated cumulative loss of each action, resulting in simple and efficient algorithms for various combinatorial structures. This idea has also been extended to online convex optimization. For instance, (hazan2007logarithmic, ) develop an FPL-style algorithm that achieves logarithmic regret for strictly convex payoff functions, particularly in settings where the decision point lies in the simplex. Follow-the-Regularized-Leader (FTRL) (shalev2006convex, ; shalev2012online, ) generalizes Follow-the-Leader by adding a regularization term to stabilize updates. (mcmahan2011follow, ) show that many online learning algorithms, including mirror descent, can be interpreted as instances of FTRL. In the context of online combinatorial optimization, (audibert2011minimax, ; audibert2014regret, ) study minimax regret bounds and propose FTRL-based algorithms alongside other strategies. (ito2021hybrid, ) further develop a hybrid algorithm based on FTRL.
Combinatorial Bandits Combinatorial bandits (cesa2012combinatorial, ) are a subclass of online combinatorial optimization where the decision-maker selects a subset of base arms from $\mathcal{A}_{full}$ in each round, forming a super arm. Early work on online shortest path problems with adversarial edge weights (gyorgy2007line, ; uchiya2010algorithms, ; liu2012adaptive, ) adapts the EXP3 framework (auer2002nonstochastic, ). Subsequent studies establish regret bounds for Thompson Sampling in structured settings (gopalan2014thompson, ) and propose general models with non-linear rewards and semi-bandit feedback (chen2016combinatorial, ; kveton2015tight, ). The matroid bandit model (kveton2014matroid, ) captures independence constraints via matroid optimization (whitney1992abstract, ). Later work reframes the shortest path problem as an adversarial MDP (rosenberg2019online, ) and introduces improved UCB variants with adaptive confidence estimation (zhou2024towards, ).
Epsilon net We identified two definitions of $\epsilon$-nets, with further discussion in Section 2.1. The first net of a given set, referring to (vershynin2018high, ), ensures that closed balls of radius $\epsilon$ centered at net points cover the given set. This relates to the covering number, that is the smallest possible cardinality of an $\epsilon$-net. This concept extends to function classes as a complexity measure, known as metric entropy, with further details in wainwright2019high , and has been used in russo2013eluder on the class of outcome functions. The second definition, as given in (matousek2013lectures, ), relates to importance sampling, requiring that every significant subset of a set intersects with the net. The classic $\epsilon$-net algorithm, introduced by haussler1986epsilon , operates via random sampling, yielding a net of size $\mathcal{O}(\frac{1}{\epsilon}\log\frac{1}{\epsilon})$. As noted in matousek2013lectures , the importance measure extends to general probability measures, inspiring our approach. Further works aim to reduce net size (mustafa2019computing, ; pach2011tight, ; rabani2009explicit, ) and address online settings (bhore2024online, ).
@@ 2 Subset Selection Framework
We consider the problem of selecting a small number of representative action from a large (potentially infinite but compact) action space, $\mathcal{A}_{full}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$. Define the expected outcome of playing action $a$ in bandit instance $\theta$ as
$\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{full},\theta\sim\mathcal{N}(0,I),$ (2)
where the bandit instance $\theta$ follows a multivariate normal distribution with zero mean and identity covariance matrix $I$. Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ forms a canonical Gaussian process (vershynin2018high, ). In fact, any constant-mean Gaussian process can be represented as the canonical process; see Supplementary B for details.
Let $\operatorname*{Conv}(\mathcal{A}_{full})$ denote the convex hull that consists of all convex combinations of finitely many actions in $\mathcal{A}_{full}$. The extreme points of the convex hull are those points $x\in\operatorname*{Conv}(\mathcal{A}_{full})$ for which there do not exist distinct points $a,a^{\prime}\in\operatorname*{Conv}(\mathcal{A}_{full})$ and $\lambda\in(0,1)$ such that $x=\lambda a+(1-\lambda)a^{\prime}$.
By the extreme point theorem, i.e., the optimum of the linear function on a convex set is attained at an extreme point—it follows that the optimal action for any bandit instance is always an extreme point of the full action space. Therefore, if we use all the extreme points as representatives of the full action space, the regret (as defined in Equation (1)) is zero. See Supplementary C for details.
This example highlights how a geometric approach can be used to solve the stochastic combination problem. However, even if one only needs the extreme points, the set of extreme points may still be large or even infinite, e.g, the extreme points of a Euclidean ball is infinite. To address this, we will later introduce the notions of $\epsilon$-nets. Without loss of generality, we assume $\mathcal{A}_{full}$ consists only of the extreme points of $\operatorname*{Conv}(\mathcal{A}_{full})$, as they are the only points of interest.
@@ 2.1 Epsilon Nets
There are at least two definitions of $\epsilon$-net: one from a geometric perspective (vershynin2018high, ) and another from a measure-theoretic perspective (matousek2013lectures, ). Let $\|\cdot\|_{2}$ denote the Euclidean norm.
  • A subset $\mathcal{A}\subseteq\mathcal{A}_{full}$ is called a Geometric $\epsilon$-net if, for all $a\in\mathcal{A}_{full}$, there exists $a^{\prime}\in\mathcal{A}$ such that
@@
  • Let $\mathcal{R}$ be a finite partition of the extreme points into disjoint clusters such that $\cup_{r\in\mathcal{R}}\;r=\mathcal{A}_{full}$. Given a measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. A subset $\mathcal{A}\subseteq\mathcal{A}_{full}$ is called a Measure-Theoretic $\epsilon$-net with respect to measure $q$ if, for any cluster $r\in\mathcal{R}$, we have:
@@
Intuitively, the geometric definition ensures point-wise closeness, while the measure-theoretic definition ensures representative sampling. In well-behaved spaces, the two coincide. For example, let $\mathcal{A}_{full}=[0,1]$ and $q$ be the uniform distribution over this interval. A grid with spacing $\epsilon$ forms a geometric $\epsilon$-net and also ensures any subinterval of measure at least $\epsilon$ contains at least a grid point.
Both definitions are useful. A geometric $\epsilon$-net ensures small regret, as it guarantees that an action close to the optimal one always exists in $\mathcal{A}$. However, it suffers from the curse of dimensionality: for $[0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$, exponentially in $n$. To address this, we incorporate the measure-theoretic $\epsilon$-net to restrict the geometric construction to only the important clusters—as determined by the $q$-measure.
The partition $\mathcal{R}$ is the central concept that bridges the two definitions of $\epsilon$-nets. If two actions $a,a^{\prime}$ are in the same cluster, they are close in the Euclidean sense, and the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., in expected squared difference). This holds because, for the canonical Gaussian process, the increments of the process correspond to the Euclidean distance:
$\|\mu_{a}-\mu_{a^{\prime}}\|_{L^{2}}=\left(\mathbb{E}(\mu_{a}-\mu_{a^{\prime}})^{2}\right)^{1/2}=\left(\mathbb{E}(a-a^{\prime})^{T}\theta\theta^{T}(a-a^{\prime})\right)^{1/2}=\|a-a^{\prime}\|_{2},\;\forall a,a^{\prime}\in\mathcal{A}_{full},$
where $\theta^{T}$ denotes the transpose of $\theta$. The middle & right equalities use Equation (2) and $\mathbb{E}\theta\theta^{T}=I$. Then, selecting one point from each cluster $r\in\mathcal{R}$ forms a geometric $\epsilon$-net (which will later be formally defined as a reference subset), where $\epsilon$ is the maximum diameter of all the clusters. This construction yields bounded expected regret, as each selected point serves as a representative for its cluster, and all other points within the same cluster have bounded deviation from the representative. Furthermore, clusters in $\mathcal{A}_{full}$ that are deemed unimportant for expected regret can be ignored: even if no representative points are selected from these clusters, the resulting expected regret will not increase significantly compared to that of the reference subset.

### v2

# Representative Action Selection for Large Action Space Meta-Bandits

@@ 1 Introduction
Different from prior approaches in multi-armed bandits (MAB) (Dani et al., 2008) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits. This problem can be seen as a large-scale combinatorial optimization under uncertainty, with applications where decisions involve a vast number of possibilities but are constrained by computational or time limitations for evaluating all options, e.g., inventory management, online recommendations.
Consider the following MAB setting: In a bandit, if a decision-maker plays an action with a fixed but unknown feature vector $a\in\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$. We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle$ where this bandit instance $\theta\in\mathbb{R}^{n}$ is drawn from an unknown multivariate Gaussian distribution. Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ forms a Gaussian process (GP) (Vershynin, 2018, Chapter 7), and we also provide results under general sub-Gaussian assumption. Our underlying setting aligns with the one considered in contextual bandit (Dani et al., 2008); however, the key difference lies in the objective.
Now consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$. For a given bandit instance $\theta$, if the best action over the full space lies within $\mathcal{A}$, the decision-maker benefits from reduced suboptimal actions to explore. Conversely, if the best action lies outside $\mathcal{A}$, regret arises from being unable to select this best action. Thus, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
$\boxed{\textnormal{Regret}:=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\;\mu_{a^{\prime}},}$ (1)
@@
This objective is motivated by practical considerations. The classic Bayesian regret in the bandit literature (Agrawal and Goyal, 2012) typically scales with the number of available actions. However, if a subset $\mathcal{A}$ is carefully chosen, the resulting Bayesian bandit regret can be significantly lower due to the reduced number of suboptimal actions. This is especially beneficial when the action space is large, as even the initialization phase can be computationally expensive. To see this, we can decompose the bandit regret as follows:
[…]
$=\mathbb{E}\sum_{t=1}^{N}\left[\max_{a\in\mathcal{A}}\mu_{a}(\theta)\!-\!\mu_{A_{t}}+\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)\!-\!\max_{a\in\mathcal{A}}\mu_{a}(\theta)\right]$
$\leq C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}].$
In the first equality, $A_{t}\in\mathcal{A}$ denotes the action chosen by a policy, e.g., Thompson Sampling (Agrawal and Goyal, 2012), in round $t$, and $N$ is the number of rounds. The expectation is taken over the randomness in the distribution of bandit instances and in actions selected by the policy. The inequality follows from the well-known regret bounds for Thompson Sampling (Lattimore and Szepesvári, 2020). $|\mathcal{A}|$ denotes the cardinality of the action set and $C>0$ is a constant. Note that if the policy has access to the full action space, the Bayesian bandit regret is instead bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$.
Our main contributions are organized as follows:
  • Meta-Bandits Framework. We propose a meta-bandits framework that specifically tackles combinatorial action selection by leveraging correlations across similar actions and bandit instances. To the best of our knowledge, this is the first such framework.
  • $\epsilon$-Net Algorithm. We introduce a simple algorithm within this framework. It starts with the intuitive idea of placing a grid over the action space, then refines it using an importance-based selection mechanism.
  • Regret Analysis. We provide theoretical guarantees for both the grid and the algorithm’s output, including upper and lower bounds on expected regret, along with results under general sub-Gaussian processes. We also discuss the cost of not using a grid, which depends on the importance-structure of the action space.
  • Generalization and Empirical Validation. We extend the analysis to settings where outcome functions are sampled from a reproducing kernel Hilbert space (RKHS), and empirically compare our algorithm to Thompson Sampling (TS) and Upper Confidence Bound (UCB).
@@ 1.1 Related Works
Multi-Armed Bandits (Lai and Robbins, 1985; Auer et al., 2002a) is defined by a set of actions (arms), each deliver outcomes that are independently drawn from a fixed and unknown distribution. The decision-maker sequentially selects an action, observes its outcome, and aims to maximize cumulative outcomes over time. Popular methods include the UCB (Auer et al., 2002a), TS (Agrawal and Goyal, 2012), and EXP3 (Auer et al., 2002b) for adversarial settings.
Optimal Action Identification focuses on identifying the action with the highest expected outcome in a MAB setting using as few independent samples as possible (Jamieson and Nowak, 2014; Kaufmann et al., 2016). Popular methods in fixed confidence setting include Action Elimination Even-Dar et al. (2006); Karnin et al. (2013), UCB, and LUCB, all of which achieve sample complexity within a $\log(|\mathcal{A}_{\operatorname*{full}}|)$ factor of the optimum. In fixed budget setting, there is Successive halving Karnin et al. (2013), successive reject Audibert and Bubeck (2010). Our algorithm includes a step for identifying the optimal action. While this line of work assumes (non-centered) sub-Gaussian outcome distributions within a bandit instance, we model each instance with a linear structure, allowing for further simplification.
Stochastic Linear Optimization assumes that the expected outcome of each action depends through the inner product between a context $\theta$ and an action $a\in\mathcal{A}_{\operatorname*{full}}$ Auer (2002); Dani et al. (2008); Rusmevichientong and Tsitsiklis (2010), and our model of each bandit instance align with this setting. This is a sequential method for maximizing a linear function—i.e., $f(a):=\mu_{a}(\theta)$ with fixed $\theta$—based on noisy observations of the function values. Soare et al. (2014) study the sample complexity of optimal action identification in this setting. However, this line of work assumes the action feature vectors are known.
GP Optimization can deal with the case that the feature vectors of actions are unknown. Srinivas et al. (2009) models the outcome function as a sample from a GP prior with a kernel function (Williams and Rasmussen, 2006). Their Bayesian algorithm, GP-UCB, uses this GP prior to first select points $x_{t}$ that enable global estimation of the function, and then plays the point with the highest posterior mean. The widespread adoption of this method in bandit settings (Valko et al., 2013; Li and Scarlett, 2022), as well as in continuous action spaces (Chowdhury and Gopalan, 2017), highlights the practicality of assuming that action outcomes are correlated. This also motivates our extension, where the outcome function of each bandit instance is modeled as a sample from a RKHS.
Before discussing two seemingly relevant lines of work that select a subset from the action space, we first highlight a advantage of our approach. These methods assume a fixed $|\mathcal{A}|=K$, which is often unclear in practice and requires restarting the algorithm when changed. In contrast, our algorithm can adapt the subset size on the fly. For example, if 100 runs all return the same action, choosing this one may suffice ($|\mathcal{A}|=1$); if each run yields a different action, we may need more iterations to ensure a small regret ($|\mathcal{A}|>100$).
[…]
Example 1.
Consider three actions: $a_{1}=[1,0]$, $a_{2}=[0.9,0.1]$, and $a_{3}=[-0.1,1]$, and suppose bandits are sampled uniformly from $\theta_{1}=[1,0]$ and $\theta_{2}=[0,1]$. Then,
$\mathbb{E}\langle a_{1},\theta\rangle=\mathbb{E}\langle a_{2},\theta\rangle=0.5,\quad\mathbb{E}\langle a_{3},\theta\rangle=0.45.$
So under the Best-222-Action perspective, $a_{1}$ and $a_{2}$ would be selected. However, this is suboptimal in our framework, since $\mu_{a_{1}}$ and $\mu_{a_{2}}$ are positively correlated:
$\mathbb{E}\max_{a\in\{a_{1},a_{2}\}}\langle a,\theta\rangle=0.55,\quad\mathbb{E}\max_{a\in\{a_{1},a_{3}\}}\langle a,\theta\rangle=1.$
In contrast, our algorithm, if run until it selects two distinct actions, would output $a_{1}$ and $a_{3}$, which is the true optimal.
Combinatorial Bandits considers that the decision maker selects $K$ of base arms from $\mathcal{A}_{\operatorname*{full}}$ in each round, forming a super arm $\mathcal{A}$, with $|\mathcal{A}|=K$. Popular methods include CUCB Chen et al. (2016), CTS Wang and Chen (2018). We argue that this line of work is not applicable to our framework: (1) It assumes that the expected outcome of a super arm depends only on the expected outcomes of its individual base arms, or imposes a stricter monotonicity condition. In our case, even though $\mathbb{E}[\mu_{a}]=c$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, super arm expected outcomes $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ can differ significantly due to correlations among actions. (2) It assumes independence across base arms, whereas we explicitly model correlations. Ignoring these correlations misses the core challenge—an issue illustrated in Example 1.
[…]
Expected supremum of Gaussian process for a given set $\mathcal{S}$ refers to the term $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]$. It is an important topic in high-dimensional probability (Vershynin, 2018). The sharpest known bounds are due to Talagrand (2014).
@@ 2 Subset Selection Framework
We consider the problem of selecting a small number of representative actions from a large action space $\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, where $n\in\mathbb{N}$. (This framework applies to the case $n=+\infty$, with the additional assumption $\sum_{i\geq 1}a^{2}_{i}\!<\!\infty$.) The expected outcome depends on both the chosen action $a\in\mathbb{R}^{n}$ and an observed context $g\in\mathbb{R}^{n}$, and includes a constant $c\in\mathbb{R}$:
$\mu_{a}:=\langle a,g\rangle+c,\quad\forall,a\in\mathcal{A}_{\operatorname*{full}},;g\sim\mathcal{N}(0,\Sigma),$ (2)
where $\Sigma$ is a positive semi-definite matrix. Let $\theta\in\mathbb{R}^{n}$ follows a multivariate normal distribution with zero mean and identity covariance matrix. The distribution of $g$ is in fact equivalent to $\Sigma^{1/2}\theta$. Now, let $\sigma_{j}$ denote the $j$-row of the matrix $\Sigma^{1/2}$. We have $g=(\langle\sigma_{j},\theta\rangle)_{j\leq n}$ and
$\langle a,g\rangle=\sum_{j\leq n}a_{j}\langle\sigma_{j},\theta\rangle=\left\langle\sum_{j\leq n}a_{j}\sigma_{j},\theta\right\rangle=\langle\Sigma^{1/2}a,\theta\rangle.$
Therefore, the setting in Equation (2) is equivalent to
$\mu_{a}:=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}},\;\theta\sim\mathcal{N}(0,I),$
where $\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}}$ denotes the image of $\mathcal{A}_{\operatorname*{full}}$ under the linear transformation $\Sigma^{1/2}$. Since the constant $c$ does not affect the regret (as defined in Equation (1)), we can, without loss of generality, focus on this canonical Gaussian process (Vershynin, 2018, Chapter 7) in the remainder:
$\boxed{\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{\operatorname*{full}},\theta\sim\mathcal{N}(0,I).}$ (3)
We define the extreme points as those $x\in\mathcal{A}_{\operatorname*{full}}$ for which there do not exist distinct $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ and $\lambda\in(0,1)$ such that $x=\lambda a+(1-\lambda)a^{\prime}$. By the extreme point theorem, if we select all extreme points—denoted $\mathcal{A}=\{a_{1},\dots,a_{K}\}$—as representatives of the full action space, the regret is zero. This is because any $a\in\mathcal{A}_{\operatorname*{full}}$ can be expressed as a convex combination of the extreme points: $a=\lambda_{1}a_{1}+\dots+\lambda_{K}a_{K}$, where $\lambda_{i}\geq 0$ and $\sum_{i=1}^{K}\lambda_{i}=1$. Thus, for any $\theta\in\mathbb{R}^{n}$:
$\begin{split}\langle a,\theta\rangle=\sum_{i=1}^{K}\lambda_{i}\langle a_{i},\theta\rangle\leq\sum_{i=1}^{K}\lambda_{i}\max_{a^{\prime}\in\mathcal{A}}\langle a^{\prime},\theta\rangle=\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}.\end{split}$
By Equation (3), it yields
$\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$
where the left inequality uses $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$. Thus, the two quantities $\max_{a\in\mathcal{A}}\mu_{a}$ and $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$ are equal.
This example highlights how a geometric approach can be used to solve the stochastic combination problem. However, even if one only needs the extreme points, the set of extreme points may still be large, e.g, the extreme points of a Euclidean ball is infinite. To address this, we will later introduce the notions of $\epsilon$-nets. Without loss of generality, we assume $\mathcal{A}_{\operatorname*{full}}$ consists only of the extreme points of $\mathcal{A}_{\operatorname*{full}}$, as they are the only points of interest.
@@ 2.1 Epsilon Nets
If the set of extreme points $\mathcal{A}_{\operatorname*{full}}$ is still large, a natural approach is to construct a grid over the action space, where the grid points serve as representative actions. This ensures that for every action in the full space, there exists a representative that is close to it. This idea is formally captured by the notion of a (geometric) $\epsilon$-net.
To proceed, we clarify what we mean by an $\epsilon$-net, as there are at least two definitions: one from a geometric perspective (Vershynin, 2018, Chapter 4) and another from a measure-theoretic perspective (Matousek, 2013, Chapter 10). Let $\|\cdot\|_{2}$ denote the Euclidean norm. Define the diameter of a compact set $r\in\mathbb{R}^{n}$ as $\mathrm{diam}(r):=\max_{a,b\in r}\|a-b\|_{2}$.
  • A subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ is called a Geometric $\epsilon$-net if, for all $a\in\mathcal{A}_{\operatorname*{full}}$, there exists $a^{\prime}\in\mathcal{A}$ such that
@@
  • Let $\mathcal{R}$ be a finite partition of the extreme points into disjoint clusters such that $\cup_{r\in\mathcal{R}}\;r=\mathcal{A}_{\operatorname*{full}}$. Given a measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. A subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ is called a Measure-Theoretic $\epsilon$-net with respect to measure $q$ if, for any cluster $r\in\mathcal{R}$, we have:
@@
A geometric $\epsilon$-net ensures small regret because if two actions $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ are close in the Euclidean sense, then the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., their expected squared difference is small):
$\|\mu_{a}\!-\!\mu_{a^{\prime}}\|_{L^{2}}=\!\left(\mathbb{E}(a\!-\!a^{\prime})^{\top}\theta\theta^{\top}(a\!-\!a^{\prime})\right)^{1/2}\!=\|a\!-\!a^{\prime}\|_{2},$
where $\theta^{\top}$ denotes the transpose of $\theta$. The equalities use Equation (3) and $\mathbb{E}\theta\theta^{\top}=I$. Therefore, by definition, a geometric $\epsilon$-net guarantees the existence of an action $a\in\mathcal{A}$ whose expected outcome $\mu_{a}$ is close to that of the optimal action for any given bandit instance. However, this net suffers from the curse of dimension: e.g., for $[0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$.
The measure-theoretic $\epsilon$-net addresses this issue. Put simply, the measure-theoretic $\epsilon$-net restricts the grid construction to only the most important clusters $r\in\mathcal{R}$, as determined by the $q$-measure.